OpenStudy (anonymous):

Help please can't figure this out!! Which value for x makes the open sentence true? 10 • 8 – 3x + x( to the 2 power)≥ x( to the 3 power)+ x( to the 2 power) A.6 B.4 C.5 D.7

OpenStudy (opcode):

Ekk you question is messy. \[ 10\times 8 – 3x + x^2≥ x^3+ x^2 \] Okay there much better. Now lets break down this question. This symbol: '≥' means greater than or equal to. So: \[(10\times 8 – 3x + x^2)≥(x^3+ x^2)\] This \(10\times 8 – 3x + x^2\) side should be greater than or equal to this side \(x^3+ x^2\). Plug in the values from your multiple choice answers till you find the value that fits the statement correctly.